# CT-UIO: Continuous-Time UWB-Inertial-Odometer Localization Using Non-Uniform B-spline with Fewer Anchors

### v1

@@ Abstract
Ultra-wideband (UWB) based positioning with fewer anchors has attracted significant research interest in recent years, especially under energy-constrained conditions. However, most existing methods rely on discrete-time representations and smoothness priors to infer a robot’s motion states, which often struggle with ensuring multi-sensor data synchronization. In this paper, we present an efficient UWB-Inertial-odometer localization system, utilizing a non-uniform B-spline framework with fewer anchors. Unlike traditional uniform B-spline-based continuous-time methods, we introduce an adaptive knot-span adjustment strategy for non-uniform continuous-time trajectory representation. This is accomplished by adjusting control points dynamically based on movement speed. To enable efficient fusion of IMU and odometer data, we propose an improved Extended Kalman Filter (EKF) with innovation-based adaptive estimation to provide short-term accurate motion prior. Furthermore, to address the challenge of achieving a fully observable UWB localization system under few-anchor conditions, the Virtual Anchor (VA) generation method based on multiple hypotheses is proposed. At the backend, we propose a CT-UIO factor graph with an adaptive sliding window for global trajectory estimation. Comprehensive experiments conducted on corridor and exhibition hall datasets validate the proposed system’s high precision and robust performance. The codebase and datasets of this work will be open-sourced at https://github.com/JasonSun623/CT-UIO.
Index Terms:
UWB-Inertial-Odometer Fusion, Continuous-Time Trajectory, Factor Graph, Few Anchors.
@@ I Introduction
Location-based services [1, 2, 3, 4] such as trajectory prediction, object tracking, and automatic picking operations, have become an integral part of robotics research. Global Navigation Satellite Systems (GNSS) can achieve meter-level accuracy in outdoor clear sky environments. However, reliable localization of GNSS is impractical indoors due to the attenuation of GNSS signals. To achieve robust indoor positioning services in GPS-denied environments, Ultra-wideband (UWB) technology is considered as a promising alternative solution benefiting from its stability, low-cost, and more scalable in large-scale deployment [5, 6]. Based on UWB signal ranging, most of the current range-based localization schemes require at least three anchors to determine a moving target’s 2D position, which is also referred to as multilateration [7, 8]. However, in energy-constrained UWB networks or deployment scenarios with spatial limitations, such as narrow corridors or tunnels, only one or two anchors may be available. This limitation renders traditional multilateration positioning methods ineffective in such environments.
@@
In such situations, range-based localization with few anchors is challenging and has become an attractive research direction. Existing methods [2, 9, 10] usually require special anchor’s antenna arrays or integrate additional sensors, such as Inertial Measurement Unit (IMU), wheel odometers, and cameras, to provide supplementary velocity or angle measurements. IMU and odometers are environment-independent and provide instantaneous response speeds, enabling the construction of relative position constraints between consecutive UWB ranging measurements. Conventional sensor fusion schemes [2, 11, 12] such as Extended Kalman Filter (EKF) and Particle Filter (PF), typically rely on smoothness priors to infer a robot’s motion states. However, these methods are based on Discrete-Time (DT) trajectory estimation, which may fail to adequately represent smoothness. Interpolation schemes used to infer robot motion between discrete states may be inaccurate. Furthermore, multi-modal sensor setups often involve asynchronous measurements from different sensors, which causes difficulty in fusing data at the same time instants during the estimation process.
@@
Within recent years, continuous-time representation has been increasingly applied to multi-sensors calibration, motion planning, and target tracking [13, 14, 15, 16]. The key advantage of continuous-time formulation lies in its ability to generate smooth trajectories with continuous data streams. In particular, B-spline defines temporal polynomials to represent trajectories by a set of control points, which enables pose querying at any time instant with locality. Most existing methods [17, 18, 19] adopt uniform the B-spline-based strategy for continuous-time trajectory smooth updates, where control points are spaced uniformly and typically predetermined. However, these methods often rely on the assumption of a constant velocity model, which is unsuitable for many real-world applications. Since the robot’s velocity is constantly changing, smaller control point spacing may result in overfitting and increased computational complexity, while larger spacing reduces overall accuracy. Benefiting from dynamic control point distribution, non-uniform B-spline is becoming an attractive alternative. Related works by Lang et al. [14] and Dubé et al. [20] illustrate that non-uniform knot sampling strategies can accommodate trajectory complexity more effectively. These methods have also proved that the non-uniform sampling strategy can be superior to the uniform strategy in terms of trajectory estimation accuracy and time performance.
@@
However, research on non-uniform B-spline methods has primarily focused on LiDAR [14] and cameras [21] applications for SLAM, it has not been applied to few-anchor UWB-based positioning. Moreover, existing non-uniform strategies for determining appropriate knot spans often rely solely on velocity estimates from IMU measurements, which are prone to noise, resulting in inaccurate spacing interval divisions.
@@
In this work, we propose CT-UIO, a continuous-time UWB-inertial-Odometer system using non-uniform B-splines with few anchors. CT-UIO dynamically adjusts control point spacing based on changes in motion trajectories. In general, although continuous-time methods have been extensively studied, most existing works rely on uniform B-splines to achieve continuous-time representations and estimation. To the best of the authors’ knowledge, this paper is the first to utilize non-uniform continuous-time representation for UWB/IMU/odometer fusion localization, we can achieve efficient pose estimation in few-anchor scenes or with fast motions through extensive experiments.
The contributions of this paper are summarized as follows.
  • We propose a continuous-time UWB-Inertial-Odometer (CT-UIO) localization system that adopts an adaptive knot span adjustment strategy for non-uniform continuous-time trajectory representation. CT-UIO fuses time-unsynchronized UWB ranging, IMU, and odometer data, enabling accurate pose estimation even in few-anchor scenes and during time-varying motions.
@@
  • We construct an IMU/odometer fusion model using an improved Extended Kalman Filter (EKF) with innovation-based adaptive estimation, providing short-term accurate motion priors. Leveraging this, the adaptive knot span adjustment strategy dynamically adjusts the number of control points based on movement speed.
@@
  • We generate virtual anchors by combining UWB ranging with motion priors from the IMU/odometer fusion model, constructing a fully observable UWB localization system under few-anchor conditions. Additionally, we propose a virtual anchors deployment scheme based on multiple hypotheses to avoid collinearity.
@@
II Related Work
@@ II-A Few-Anchor UWB Localization
Few-anchor UWB localization has recently attracted increasing attention and has been widely studied due to its low complexity and cost-effective deployment [9, 10, 22]. These research works focus exclusively on fusion schemes that integrate UWB measurements with other motion sensors. Mainstream works primarily employ filter-based, optimization-based, and two-stage methods to incorporate additional velocity observations. In filter-based solutions, Cao et al. [23] employed Extended Kalman Filtering (EKF) to combine a single UWB anchor with a nine-DOF IMU for velocity and orientation estimation. Qin et al. [24] designed a Set Membership Filter (SMF) method based on constrained zonotopes for single-beacon localization problems. Gao et al. [22] introduced UWB ranging, non-holonomic constraints from the carrier, and trajectory constraints as observations of the EKF system. The proposed method can realize positioning in corridor-like areas. However, it relies on non-holonomic and corridor-specific trajectory constraints, which may not be applicable in unknown environmental structures. In optimization-based solutions, Li et al. [25] proposed a Moving Horizon Estimation (MHE) algorithm, which utilizes historical measurement data from a single UWB anchor and IMU during a time horizon. Additionally, a Gradient Aware Levenberg-Marquardt (GALM) algorithm was further proposed to solve the optimization problem with a calculation cost. For two-stage methods, Yang et al. [26] proposed an improved positioning method that leverages Bayesian optimization to provide the estimated IMU data based on an error model. A filter then updates the position at the ranging time using the prior position error. Zhou et al. [27] proposed a two-stage optimization method that could yield accurate solutions with the single-anchor and odometer. In the first stage, they construct a factor graph that incorporates odometry positions and UWB measurements to optimize state estimation. In the second stage, an adaptive trust region algorithm is performed to refine the location estimate and maintain robustness with inequality constraints. While these methods approximate the time offsets between state estimation and sensor measurements, the hypothesis of time synchronization may not always hold in practical scenarios. Unlike the existing few-anchor UWB localization methods, our method formulates the trajectory in continuous-time with B-spline, which allows asynchronous, high-frequency sensor measurements to be aligned at any time instants along the trajectory, rather than relying on linear interpolation of discrete-time poses only at measurement times.
@@ II-B Continuous-Time Representation Using B-spline
Continuous-time representation is a popular and natural choice for formulating trajectories and ensuring smoothness. It can estimate the robot’s pose as a continuous function of time without the need to introduce additional states at every measurement time. The most common continuous-time-based model is based on splines and Gaussian processes [15]. In this article, we focus on B-spline-based trajectory representation, current literature has paid increasing attention to applying B-splines for continuous-time state estimation in multi-sensor fusion systems, including the LiDAR-Inertial system, visual-inertial system, ultra-wideband-inertial and so on. Nguyen et al. [16] developed a real-time continuous-time LiDAR-inertial odometry (SLICT2), which achieved efficient optimization with few iterations using a simple solver. Lu et al. [28] proposed an event-based visual-inertial velometer that incrementally incorporates measurements from a stereo event camera and IMU. Li et al. [17] proposed a spline-based approach (SFUISE) for continuous-time Ultra-wideband-Inertial sensor fusion, which addressed the limitations of discrete-time sensor fusion schemes in asynchronous multi-sensor fusion and online calibration. These methods generally employ uniform knot B-splines for trajectory modeling, relying on assumptions of zero velocity or constant speed. However, these assumptions often fail to capture the dynamic nature of real-world motion. Non-uniform B-splines offer a more flexible distribution of control points, enabling a different density of control points based on the smoothness of the trajectory segment. Ovrén et al. [29] introduced an energy proportionality index in spline fitting to optimize the selection of knot spacing. By leveraging the different frequency response characteristics of spline basis functions and specific energy values, appropriate knot spacing can be automatically selected. Lang et al. [14] tightly coupled the measurements from LiDAR, IMU, and cameras using non-uniform B-spline curves. They adaptively adjusted the number of control points based on IMU observations to detect different motion patterns, thereby improving adaptability to complex environments and motion patterns. However, relying solely on IMU data leads to velocity estimate divergence, which impacts the distribution of control points and ultimately reduces positioning accuracy. In contrast, the proposed method integrates an IMU/Odometer fusion model to provide short-term accurate velocity estimates, enabling the dynamic adjustment of control point density for more precise trajectory modeling.
@@
Figure 1: The framework of CT-UIO: In the preprocessing stage, UWB ranging data is detected, and outliers are removed. Simultaneously, incoming IMU and odometer data are fused using an adaptive EKF to provide short-term accurate motion priors. In the front-end, the results of the IMU/odometer fusion model are combined with UWB ranging to generate virtual anchors. Additionally, Based on the motion estimates from the IMU/odometer fusion model, the adaptive knot span adjustment strategy non-uniformly places control points. In the back-end, we conduct a CT-UIO factor graph with an adaptive sliding window for global trajectory estimation.
III System Overview

### v2

@@ Abstract
Ultra-wideband (UWB) based positioning with fewer anchors has attracted significant research interest in recent years, especially under energy-constrained conditions. However, most existing methods rely on discrete-time representations and smoothness priors to infer a robot’s motion states, which often struggle with ensuring multi-sensor data synchronization. In this article, we present a continuous-time UWB-Inertial-Odometer localization system (CT-UIO), utilizing a non-uniform B-spline framework with fewer anchors. Unlike traditional uniform B-spline-based continuous-time methods, we introduce an adaptive knot-span adjustment strategy for non-uniform continuous-time trajectory representation. This is accomplished by adjusting control points dynamically based on movement speed. To enable efficient fusion of inertial measurement unit (IMU) and odometer data, we propose an improved extended Kalman filter (EKF) with innovation-based adaptive estimation to provide short-term accurate motion prior. Furthermore, to address the challenge of achieving a fully observable UWB localization system under few-anchor conditions, the virtual anchor (VA) generation method based on multiple hypotheses is proposed. At the backend, we propose an adaptive sliding window strategy for global trajectory estimation. Comprehensive experiments are conducted on three self-collected datasets with different UWB anchor numbers and motion modes. The result shows that the proposed CT-UIO achieves $0.403m$, $0.150m$, and $0.189m$ localization accuracy in corridor, exhibition hall, and office environments, yielding $17.2\%$, $26.1\%$, and $15.2\%$ improvements compared with competing state-of-the-art UIO systems, respectively. The codebase and datasets of this work will be open-sourced at https://github.com/JasonSun623/CT-UIO.
@@ I Introduction
Location-based services [9184260, 10538051, 9705111, 10679703, Diff9D] such as trajectory prediction, object tracking, and automatic picking operations, have become an integral part of robotics research. Global navigation satellite systems (GNSS) can achieve meter-level accuracy in outdoor clear-sky environments. However, reliable localization of GNSS is impractical indoors due to the attenuation of GNSS signals. To achieve robust indoor positioning services in GPS-denied environments, Ultra-wideband (UWB) technology is considered as a promising alternative solution benefiting from its stability, low cost, and scalability in large-scale deployment [10400775, 10704747]. Based on UWB signal ranging, most of the current range-based localization schemes require at least three anchors to determine a moving target’s 2D position, which is also referred to as multilateration [2013.0815, PARK2024103234]. Compared to the range-based localization schemes, the fingerprint estimation (FPE) could provide relatively high localization accuracy under multipath and non-line-of-sight (NLOS) scenarios [8806041]. In recent years, deep-learning-based UWB localization has emerged as a rapidly evolving field. Poulose and Han [app10186290] proposed a long short-term memory (LSTM) based model to mitigate positioning degradation caused by UWB ranging errors and NLOS conditions, which achieved improved performance with a mean localization error of $7cm$. However, in energy-constrained UWB networks or deployment scenarios with spatial limitations, such as narrow corridors or tunnels, only one or two anchors may be available. These limitations render traditional multilateration, fingerprinting, and even deep-learning-based methods ineffective in such environments.
@@
In such situations, range-based localization with few anchors is challenging and has become an attractive research direction. Existing methods [10268655, tong2022single, 9829196, 9440892] typically require special anchors’ antenna arrays or integrate additional sensors, such as inertial measurement unit (IMU), wheel odometers, and cameras, to provide supplementary velocity or angle measurements. IMU and odometers are environment-independent and provide instantaneous response speeds, enabling the construction of relative position constraints between consecutive UWB ranging measurements. Conventional sensor fusion schemes [10268655, 9440987, Xiong2022] such as extended Kalman filter (EKF) and particle filter (PF), typically rely on smoothness priors to infer a robot’s motion states. However, these methods are based on discrete-time (DT) trajectory estimation, which may fail to adequately represent smoothness. Interpolation schemes used to infer robot motion between discrete states may be inaccurate. Furthermore, multi-modal sensor setups often involve asynchronous measurements from different sensors, which causes difficulty in fusing data at the same time instants during the estimation process.
@@
Within recent years, continuous-time (CT) representation has been increasingly applied to multi-sensor calibration, motion planning, and target tracking [10311080, lang2023coco, cioffi2022continuous, nguyen2024eigen]. The key advantage of the continuous-time formulation lies in its ability to generate smooth trajectories with continuous data streams. In particular, B-spline defines temporal polynomials to represent trajectories by a set of control points, which enables pose querying at any time instant with locality. Most existing methods [li2023continuous, 10045587, 9636676] adopt the uniform B-spline-based strategy for continuous-time trajectory smooth updates, where control points are spaced uniformly and typically predetermined. However, these methods often rely on the assumption of a constant velocity model, which is unsuitable for many real-world applications. Since the robot’s velocity is constantly changing, smaller control point spacing may result in overfitting and increased computational complexity, while larger spacing reduces overall accuracy. Benefiting from dynamic control point distribution, non-uniform B-splines are becoming an attractive alternative. Related research works by Lang et al. [lang2023coco] and Dubé et al. [dube2016non] illustrate that non-uniform knot sampling strategies can accommodate trajectory complexity more effectively. These methods have also proved that the non-uniform sampling strategy can be superior to the uniform strategy in terms of trajectory estimation accuracy and time performance.
@@
However, research on non-uniform B-spline methods has primarily focused on light detection and ranging (LiDAR) [lang2023coco] and cameras [9320417] applications for simultaneous localization and mapping (SLAM), it has not been applied to few-anchor UWB-based positioning. Moreover, existing non-uniform strategies for determining appropriate knot spans often rely solely on velocity estimates from IMU measurements, which are prone to noise, resulting in inaccurate spacing interval divisions.
@@
In this work, we propose a continuous-time UWB-Inertial-Odometer system (CT-UIO) using non-uniform B-splines with few anchors. CT-UIO dynamically adjusts control point spacing based on changes in motion trajectories. In general, although continuous-time methods have been extensively studied, most existing works rely on uniform B-splines to achieve continuous-time representations and estimation. To the best of the authors’ knowledge, this article is the first to utilize a non-uniform continuous-time representation for UWB/IMU/odometer fusion localization, which can achieve efficient pose estimation in few-anchor scenes or with time-varying motion states.
The contributions of this article are summarized as follows.
  • We propose the first continuous-time UWB-Inertial-Odometer localization system (CT-UIO) that employs a non-uniform B-spline trajectory representation. This enables seamless fusion of time-unsynchronized UWB, IMU, and odometer measurements, achieving both high positioning accuracy and real-time performance in few-anchor, time-varying motion conditions.
@@
  • We develop an improved EKF-based IMU/odometer fusion model with innovation-based adaptive estimation, providing accurate short-term motion priors. These priors are then used to dynamically adjust knot spans in the trajectory optimization, increasing control points only when motion speed changes significantly. This adaptive knot span adjustment strategy allows accurate modeling of rapid motion changes and improves overall pose estimation accuracy.
@@
  • We propose a virtual anchor (VA) generation method that combines UWB ranging with motion priors from the improved EKF-based IMU/odometer fusion model, ensuring a fully observable UWB localization system even when fewer physical UWB anchors are available. Additionally, we design a multiple-hypothesis-based VA deployment scheme to avoid collinearity issues and enhance positioning robustness.
@@
The remainder of this article is organized as follows. In Sec. II, the related works for few-anchor UWB localization and continuous-time representation are discussed. In Sec. III, the motivation is presented. In Sec. IV, the proposed CT-UIO is introduced in detail. To verify the performance of the proposed methods, extensive experiments are conducted in Sec. V. Finally, in Sec. VI, we draw the conclusion and discuss future research directions.
II Related Works
@@ II-A Few-Anchor UWB Localization
Few-anchor UWB localization has recently attracted increasing attention and has been widely studied due to its low complexity and cost-effective deployment [tong2022single, cao2020accurate, qin2024single]. These research works focus exclusively on fusion schemes that integrate UWB measurements with other motion sensors. Mainstream works primarily employ filter-based, optimization-based, and two-stage methods to incorporate additional velocity observations. In filter-based solutions, Cao et al. [cao2020accurate] employed EKF to combine a single UWB anchor with a nine-DOF IMU for velocity and orientation estimation. Qin et al. [qin2024single] designed a set membership filter (SMF) method based on constrained zonotopes for single-beacon localization problems. Gao et al. [penggang2022novel] introduced UWB ranging, non-holonomic constraints from the carrier, and trajectory constraints as observations of the EKF system. The proposed method can realize positioning in corridor-like areas. However, it relies on non-holonomic and corridor-specific trajectory constraints, which may not be applicable in unknown environmental structures. In optimization-based solutions, Dong et al. [9829196] proposed a velocity-constrained Moving Horizon Estimation (VMHE). By formulating the trajectory optimization problem through incorporating indirectly derived velocity constraints and the geometric evolution equation. Experimental results demonstrated that the proposed method achieves robust trajectory estimation in long-duration flights using only a single UWB beacon. Li et al. [li2021computationally] proposed a Moving Horizon Estimation (MHE) algorithm, which utilizes historical measurement data from a single UWB anchor and IMU during a time horizon. Additionally, a Gradient Aware Levenberg-Marquardt (GALM) algorithm was further proposed to solve the optimization problem with a calculation cost. For two-stage methods, Yang et al. [yang2023novel] proposed an improved positioning method that leverages Bayesian optimization to provide the estimated IMU data based on an error model. A filter then updates the position at the ranging time using the prior position error. Zhou et al. [zhou2024optimization] proposed a two-stage optimization method that could yield accurate solutions with the single-anchor and odometer. In the first stage, they construct a factor graph that incorporates odometry positions and UWB measurements to optimize state estimation. In the second stage, an adaptive trust region algorithm is performed to refine the location estimate and maintain robustness with inequality constraints. Cheng et al. [WOS:001173317800046] proposed the loosely coupled fusion framework utilizing generalized probability data association (GPDA). This method effectively associates UWB measurements with INS-derived state estimates, reducing the impact of NLOS error. While these methods approximate the time offsets between state estimation and sensor measurements, the hypothesis of time synchronization may not always hold in practical scenarios. Unlike the existing few-anchor UWB localization methods, our method formulates the trajectory in continuous-time with B-spline, which allows asynchronous, high-frequency sensor measurements to be aligned at any time instants along the trajectory, rather than relying on linear interpolation of discrete-time poses only at measurement times.
@@ II-B Continuous-Time Representation Using B-spline
Continuous-time (CT) representation is a popular and natural choice for formulating trajectories and ensuring smoothness. It can estimate the robot’s pose as a continuous function of time without the need to introduce additional states at every measurement time. The most common continuous-time-based model is based on splines and Gaussian processes [cioffi2022continuous]. In this article, we focus on B-spline-based trajectory representation. Current literature has paid increasing attention to applying B-splines for continuous-time state estimation in multi-sensor fusion systems, including the LiDAR-Inertial system, visual-inertial system, ultra-wideband-inertial, and so on. Nguyen et al. [nguyen2024eigen] developed a real-time continuous-time LiDAR-inertial odometry (SLICT2), which achieved efficient optimization with few iterations using a simple solver. Lu et al. [lu2023event] proposed an event-based visual-inertial velometer that incrementally incorporates measurements from a stereo event camera and IMU. Li et al. [li2023continuous] proposed a spline-based approach (SFUISE) for continuous-time Ultra-wideband-Inertial sensor fusion, which addressed the limitations of discrete-time sensor fusion schemes in asynchronous multi-sensor fusion and online calibration. These methods generally employ uniform knot B-splines for trajectory modeling, relying on assumptions of zero velocity or constant speed. However, these assumptions often fail to capture the dynamic nature of real-world motion. Non-uniform B-splines offer a more flexible distribution of control points, enabling a different density of control points based on the smoothness of the trajectory segment. Ovrén et al. [ovren2018spline] introduced an energy proportionality index in spline fitting to optimize the selection of knot spacing. By leveraging the different frequency response characteristics of spline basis functions and specific energy values, appropriate knot spacing can be automatically selected. Lang et al. [lang2023coco] tightly coupled the measurements from LiDAR, IMU, and cameras using non-uniform B-spline curves. They adaptively adjusted the number of control points based on IMU observations to detect different motion patterns, thereby improving adaptability to complex environments and motion patterns. However, relying solely on IMU data leads to velocity estimate divergence, which impacts the distribution of control points and ultimately reduces positioning accuracy. In contrast, the proposed method integrates an IMU/Odometer fusion model to provide short-term accurate velocity estimates, enabling the dynamic adjustment of control point density for more precise trajectory modeling.
((a))
((b))
@@
Figure 1: The framework of CT-UIO: In the preprocessing stage, UWB ranging data is detected, and outliers are removed. Simultaneously, incoming IMU and odometer data are fused using an improved EKF to provide short-term accurate motion priors. The frontend performs incremental pose updates of the robot using the IMU/odometer fusion model, and the resulting poses are combined with UWB ranging data to determine VAs’ positions. Furthermore, leveraging the motion estimates from the IMU/odometer fusion model, the adaptive knot span adjustment strategy non-uniformly places the control points. The backend conducts factor graph optimization with an adaptive sliding window to achieve global trajectory estimation.
III Motivation
Despite the progress in few-anchor UWB localization and continuous-time trajectory estimation, these methods still suffer from several challenges related to sensor synchronization, adaptability, and observability under few-anchor conditions. The primary challenges can be summarized as follows.
Observability degradation and asynchronous fusion: When only one or two UWB anchors are available, as commonly encountered in corridors or energy-constrained deployments. Traditional multilateration or range-based optimization methods become under-constrained due to the lack of sufficient observations. Besides, discrete-time fusion frameworks estimate motion states only at discrete sampling instants. This formulation depends on interpolation and smoothness priors to approximate trajectories, leading to motion distortion and drift when fusing asynchronous UWB, IMU, and odometer data.
Limitations of uniform continuous-time representations: Most existing CT methods employ uniform B-spline representations with fixed knot intervals, assuming approximately constant motion. However, such uniform spacing cannot adapt to time-varying motion, causing overfitting and increased computational cost during slow motion, and loss of accuracy during rapid motion.
Motivated by the above discussion, we propose the Continuous-Time UWB-Inertial-Odometer (CT-UIO) framework that introduces an adaptive knot-span adjustment strategy for non-uniform B-spline trajectory modeling and employs an improved EKF-based IMU/odometer fusion model to generate virtual anchors (VAs), thereby ensuring full observability under few-anchor conditions.
IV System Overview